OpenStudy (campbell_st):

how many different shaped rectangles, with positive integer dimensions, have a perimeter equal to their area?

OpenStudy (campbell_st):

and what are the dimensions

OpenStudy (anonymous):

So this? \[ nk = 2(n+k) \]

OpenStudy (campbell_st):

yes

OpenStudy (anonymous):

Are there any other restrictions on \(n\) and \(k\)? Seems like there are infinite solutions.

OpenStudy (campbell_st):

nope... only positive integers

OpenStudy (campbell_st):

thought I was a nice question of the day and it's high school maths

OpenStudy (freckles):

k cannot be 2 if we write n as a function of k

OpenStudy (misty1212):

i got \(6\) and \(3\)

OpenStudy (freckles):

\[nk=2(n+k) \\ nk-2n=2k \\ n(k-2)=2k \\ n=\frac{2k}{k-2} \\ k \neq 2 \\ \text{ pluggin \in k for 2 we get } 2n=2(n+2) \\ 2n=2n+4 \\0 \neq 4 \]

OpenStudy (cwrw238):

3 and 6 works

OpenStudy (cwrw238):

k = 2n/ (n - 2)

OpenStudy (freckles):

right so n cannot be 2 either

OpenStudy (campbell_st):

well 6 and 3 is correct... there is 1 more

OpenStudy (misty1212):

i guess you need \(n-2|2n\)

OpenStudy (misty1212):

i meant 4 and 4

OpenStudy (cwrw238):

k=4 . n=4 works

OpenStudy (campbell_st):

correct... that's it

ganeshie8 (ganeshie8):

why only those two are the solutions ?

OpenStudy (freckles):

\[n=\frac{2k}{k-2}=2+\frac{4}{k-2}\] k-2 has to divide 4 -- 3-2 divides 4 4-2 divides 4 5-2 doesn't 6-2 doesn't 7-2 exceeds 4 so there was only 2 right?

OpenStudy (campbell_st):

because that's what the maths says \[2l + 2w = lw\] then \[w = \frac{2l}{l - 2}\] then I did this \[w = \frac{2l - 4}{l -2} + \frac{4}{l - 2}\] which gives \[w = 2 + \frac{4}{l - 2}\] so remembering the dimensions are positive integer values... the only way \[\frac{4}{l -2}\] can be a positive integer is if (l - 2) = 1, 2 or 4 so for l - 2 to be positive and an integer then l - 2 must have l = 3, 4 or 6 substitute the length values to find the widths... and so only 2 solutions hope it makes sense

ganeshie8 (ganeshie8):

Nice :)